# Spec-VLA: Speculative Decoding for Vision-Language-Action Models with Relaxed Acceptance

### v1

@@ Abstract
Vision-Language-Action (VLA) models have made substantial progress by leveraging the robust capabilities of Visual Language Models (VLMs). However, VLMs’ significant parameter size and autoregressive (AR) decoding nature impose considerable computational demands on VLA models. While Speculative Decoding (SD) has shown efficacy in accelerating Large Language Models (LLMs) by incorporating efficient drafting and parallel verification, allowing multiple tokens to be generated in one forward pass, its application to VLA models remains unexplored. This work introduces Spec-VLA, an SD framework designed to accelerate VLA models. Due to the difficulty of the action prediction task and the greedy decoding mechanism of the VLA models, the direct application of the advanced SD framework to the VLA prediction task yields a minor speed improvement. To boost the generation speed, we propose an effective mechanism to relax acceptance utilizing the relative distances represented by the action tokens of the VLA model. Empirical results across diverse test scenarios affirm the effectiveness of the Spec-VLA framework, and further analysis substantiates the impact of our proposed strategies, which enhance the acceptance length by , achieving speedup compared with the OpenVLA baseline, without compromising the success rate. The success of the Spec-VLA framework highlights the potential for broader application of speculative execution in VLA prediction scenarios.
@@ 1 Introduction
To achieve the goals above, the parameter size of backbone VLMs is substantial, increasing the computational demand for robot control systems. Meanwhile, the VLMs’ Autoregressive (AR) next-token-prediction strategy further increases the decoding latency of VLA models. A series of studies address the efficiency issue through model architecture redesign Wen et al. (2025); Liu et al. (2024) or task-specific optimizations Kim et al. (2025). Other efforts incorporate Large Language Model (LLM) inference acceleration methods such as Early-Exit Schuster et al. (2022) and Jacobi-Decoding Kou et al. (2024) into VLA inference Yue et al. (2024); Song et al. (2025). However, incorporating such methods requires resource-intensive fine-tuning of the backbone VLM for Early-Exit Yue et al. (2024) or pretraining for Jacobi-Decoding Song et al. (2025). Moreover, in Jacobi-Decoding, enabling parallel decoding degrades the model performance compared to AR decoding Song et al. (2025).
Speculative Decoding (SD) Leviathan et al. (2023) provides a lossless solution and also allows for the parallel generation of LLMs. A typical SD architecture, such as Eagle Li et al. (2024d), employs a draft model to generate draft tokens efficiently, with the LLMs serving as the verification model to ensure the correctness of these tokens. As the parameters of the draft model are decoupled, additional fine-tuning of the verification model is not required.
@@
This work introduces the speculative decoding framework to the AR robot action generation. We propose the Spec-VLA, the first SD framework designed for VLA inference acceleration, which applies the advanced features of the speculative decoding to the robot action generation scenarios. Surprisingly, the direct application of the SD framework yields minor speed improvements due to the intricate difficulty of VLA prediction for the draft model and the greedy decoding strategy. To further boost the generation speed, we propose utilizing VLA models’ token representation to relax the acceptance based on the action distance between draft tokens and ground-truth tokens. Empirical results across various test scenarios demonstrate the effectiveness of the Spec-VLA framework, enabling an acceptance length from 2.10 to 2.94. Analysis confirms that our proposed relaxation of acceptance strategy significantly enhances the acceptance length by to , enhances the generation speed by to while maintaining the success rate of the VLA models. Our findings validate the robustness of the VLA models, demonstrating the potential of SD frameworks in the VLA prediction domain.
@@ 2.1 Acceleration for VLA Models
Structural modifications, such as Robomamba Liu et al. (2024) and TinyVLA Wen et al. (2025), redesign model backbones using lightweight SSM or distilled vision encoders, achieving latency reduction through structural simplification. The Kim et al. (2025) propose temporal consistency losses to regularize action smoothness, and Song et al. (2025) reformulate decoding via Jacobi iteration for parallel trajectory generation. The aforementioned methodologies not only require domain-specific data fine-tuning or retraining but also introduce augmented system complexity through model architectural redesign.
@@ 2.2 Speculative Decoding for LLMs
The SD has emerged as an effective paradigm for inference acceleration in AR generative models, such as machine translation models Stern et al. (2018) and decoder-only LLMs Chen et al. (2023). The evolutionary trajectory of SD frameworks reveals three distinct development phases. Pioneering SD frameworks exemplified by Medusa Cai et al. (2024) and Medusa-CTC Wen et al. (2024) introduced parallel generation capabilities through multi-head decoding architectures coupled with tree-attention verification mechanisms. Subsequent developments in the Eagle series, including Eagle Zhang et al. (2024a) and Eagle-2 Li et al. (2024c), advanced the paradigm through architectural innovations in draft modeling, achieving superior speedup ratios via high-quality draft token generation. Recently, the Eagle-3 Li et al. (2025) and HASS Zhang et al. (2024b) have further improved the generation capabilities by employing a training-time testing strategy.
@@ 6.1 Acceptance Length Proportion
Table 2 quantifies the distribution of acceptance lengths (0–5) under the Spec-VLA framework, comparing non-relaxed and relaxed conditions across four datasets. The data reveals a distinct trend: non-relaxed acceptance disproportionately favors shorter sequences (lengths 0–1), with proportions sharply declining for longer lengths (2–5), whereas relaxed acceptance exhibits a more balanced distribution. The dominance of short sequences under non-relaxed conditions (e.g., 50.24% at length 0) highlights a critical inefficiency: models prioritize "safe" short predictions to avoid constraint violations. This artificially low conversion rate for longer sequences implies that strict constraints act as a bottleneck preventing the model from predicting longer action sequence. The most pronounced contrast occurs in Libero-Object at length 4: non-relaxed acceptance plummets to 0.56% versus 6.22% under relaxed conditions—an 11-fold relative increase. Similarly, Libero-Long exhibits dramatic divergence at length 4 (0.50% vs. 3.35%, 6.7× improvement) and Libero-Spatial at length 3 (2.46% vs. 8.23%, 3.3× improvement). Even at maximum length 5, relaxed acceptance achieves non-zero proportions (e.g., 0.53% for relaxed in Libero-Object vs. 0% non-relaxed). These disparities highlight a critical limitation of strict constraints: they disproportionately penalize longer sequences. Relaxation alleviates this by allowing semantically compatible draft tokens to be accepted, thereby increasing sequence diversity without compromising task success rates.

### v2

@@ Abstract
Vision-Language-Action (VLA) models have made substantial progress by leveraging the robust capabilities of Visual Language Models (VLMs). However, VLMs’ significant parameter size and autoregressive (AR) decoding nature impose considerable computational demands on VLA models. While Speculative Decoding (SD) has shown efficacy in accelerating Large Language Models (LLMs) by incorporating efficient drafting and parallel verification, allowing multiple tokens to be generated in one forward pass, its application to VLA models remains unexplored. This work introduces Spec-VLA, an SD framework designed to accelerate VLA models. Due to the difficulty of the action prediction task and the greedy decoding mechanism of the VLA models, the direct application of the advanced SD framework to the VLA prediction task yields a minor speed improvement. To boost the generation speed, we propose an effective mechanism to relax acceptance utilizing the relative distances represented by the action tokens of the VLA model. Empirical results across diverse test scenarios affirm the effectiveness of the Spec-VLA framework, and further analysis substantiates the impact of our proposed strategies, which enhance the acceptance length by , achieving speedup compared with the OpenVLA baseline, without compromising the success rate. The success of the Spec-VLA framework highlights the potential for broader application of speculative execution in VLA prediction scenarios. We make our code and data publicly available at https://github.com/PineTreeWss/SpecVLA.
@@ 1 Introduction
To achieve the goals above, the parameter size of backbone VLMs is substantial, increasing the computational demand for robot control systems. Meanwhile, the VLMs’ Autoregressive (AR) next-token-prediction strategy further increases the decoding latency of VLA models. A series of studies address the efficiency issue through model architecture redesign Wen et al. (2025); Liu et al. (2024b) or task-specific optimizations Kim et al. (2025). Other efforts incorporate Large Language Model (LLM) inference acceleration methods such as Early-Exit Schuster et al. (2022) and Jacobi-Decoding Kou et al. (2024) into VLA inference Yue et al. (2024); Song et al. (2025). However, incorporating such methods requires resource-intensive fine-tuning of the backbone VLM for Early-Exit Yue et al. (2024) or pretraining for Jacobi-Decoding Song et al. (2025). Moreover, in Jacobi-Decoding, enabling parallel decoding degrades the model performance compared to AR decoding Song et al. (2025).
Speculative Decoding (SD) provides a lossless solution and also allows for the parallel generation of LLMs. A typical SD architecture, such as Eagle Li et al. (2024d), employs a draft model to generate draft tokens efficiently, with the LLMs serving as the verification model to ensure the correctness of these tokens. As the parameters of the draft model are decoupled, additional fine-tuning of the verification model is not required.
@@
This work introduces the speculative decoding framework to the AR robot action generation. We propose the Spec-VLA, the first SD framework designed for VLA inference acceleration, which applies the advanced features of the speculative decoding to the robot action generation scenarios. Surprisingly, the direct application of the SD framework yields minor speed improvements due to the intricate difficulty of VLA prediction for the draft model and the greedy decoding strategy. To further boost the generation speed, we propose utilizing VLA models’ token representation to relax the acceptance based on the action distance between draft tokens and ground-truth tokens. Empirical results across various test scenarios demonstrate the effectiveness of the Spec-VLA framework, enabling an acceptance length from 2.10 to 2.94. Analysis confirms that our proposed relaxation of acceptance strategy significantly enhances the acceptance length by to , enhances the generation speed by to while maintaining the success rate of the VLA models.
Existing reinforcement learning studies have highlighted the importance of robustness, either in multi-task settings where policy degradation occurs Bai et al. (2023); Liu et al. (2025), or under adversarial perturbations where agents exhibit vulnerability Bai et al. (2025). Inspired by these studies, we further analyze the robustness of VLA models under speculative decoding with relaxed acceptance by exploring the relaxation threshold under multiple tasks, demonstrating the potential of SD frameworks in the VLA prediction domain.
@@ 2.1 Acceleration for VLA Models
Structural modifications, such as Robomamba Liu et al. (2024b) and TinyVLA Wen et al. (2025), redesign model backbones using lightweight SSM or distilled vision encoders, achieving latency reduction through structural simplification. The Kim et al. (2025) propose temporal consistency losses to regularize action smoothness, and Song et al. (2025) reformulate decoding via Jacobi iteration for parallel trajectory generation. The aforementioned methodologies not only require domain-specific data fine-tuning or retraining but also introduce augmented system complexity through model architectural redesign. Beyond architectural redesign and decoding strategies, recent RL-based studies also target efficiency in VLA inference. SEER Bai et al. (2024) improves sample efficiency via aligned experience estimation and policy regularization, while D3P Yu et al. (2025) accelerates inference by adaptively adjusting diffusion steps. Together, they highlight complementary RL-driven strategies for efficient VLA inference.
@@ 2.2 Speculative Decoding for LLMs
The SD has emerged as an effective paradigm for inference acceleration in AR generative models, such as machine translation models Stern et al. (2018) and decoder-only LLMs Chen et al. (2023). The evolutionary trajectory of SD frameworks reveals three distinct development phases. Pioneering SD frameworks exemplified by Medusa Cai et al. (2024) and Medusa-CTC Wen et al. (2024) introduced parallel generation capabilities through multi-head decoding architectures coupled with tree-attention verification mechanisms. Subsequent developments in the Eagle series, including Eagle Zhang et al. (2024a) and Eagle-2 Li et al. (2024c), advanced the paradigm through architectural innovations in draft modeling, achieving superior speedup ratios via high-quality draft token generation. Recently, the Eagle-3 Li et al. (2025) and HASS Zhang et al. (2024b) have further improved the generation capabilities by employing a training-time testing strategy. The framework have shown remakable superiority for LLM acclleration ( - ), compared with Jacobi-Decoding ( - ) Kou et al. (2024) and Early-Exit Decoding ( - ) Liu et al. (2024a).
@@ 6.1 Acceptance Length Proportion
Table 2 quantifies the distribution of acceptance lengths (0–5) under the Spec-VLA framework, comparing non-relaxed and relaxed conditions across four datasets. The data reveals a distinct trend: non-relaxed acceptance disproportionately favors shorter sequences (lengths 0–1), with proportions sharply declining for longer lengths (2–5), whereas relaxed acceptance exhibits a more balanced distribution. The dominance of short sequences under non-relaxed conditions (e.g., 50.24% at length 0) highlights a critical inefficiency: models prioritize ‘safe’ short predictions to avoid constraint violations. This artificially low conversion rate for longer sequences implies that strict constraints act as a bottleneck preventing the model from predicting longer action sequence. The most pronounced contrast occurs in Libero-Object at length 4: non-relaxed acceptance plummets to 0.56% versus 6.22% under relaxed conditions—an 11-fold relative increase. Similarly, Libero-Long exhibits dramatic divergence at length 4 (0.50% vs. 3.35%, 6.7× improvement) and Libero-Spatial at length 3 (2.46% vs. 8.23%, 3.3× improvement). Even at maximum length 5, relaxed acceptance achieves non-zero proportions (e.g., 0.53% for relaxed in Libero-Object vs. 0% non-relaxed). These disparities highlight a critical limitation of strict constraints: they disproportionately penalize longer sequences. Relaxation alleviates this by allowing semantically compatible draft tokens to be accepted, thereby increasing sequence diversity without compromising task success rates.
@@ 7 Conclusion
Acknowledgements
This work was supported in part by the Science and Technology Development Fund of Macau SAR (Grant No. FDCT/0007/2024/AKP), the Science and Technology Development Fund of Macau SAR (Grant No. FDCT/0070/2022/AMJ, China Strategic Scientific and Technological Innovation Cooperation Project Grant No. 2022YFE0204900), the Science and Technology Development Fund of Macau SAR (Grant No. FDCT/060/2022/AFJ, National Natural Science Foundation of China Grant No. 62261160648), the UM and UMDF (Grant Nos. MYRG-GRG2023-00006-FST-UMDF, MYRG-GRG2024-00165-FST-UMDF, EF2024-00185-FST), and the National Natural Science Foundation of China (Grant No. 62266013).
@@ Appendix B Spec-VLA Decoding Algorithm
Appendix C Accelleration for Quantilized Models
To illustrate SD’s complementary potential, we explored further by combining Speculative Decoding with quantization.
We compared inference speeds for the OpenVLA model using int8, int4 quantization, and BF16 representations on the Tesla A100 (80G) GPU. We observed that the int8 and int4 quantization lead to decreased inference speed (Table 5). This result may be attributed to the additional overhead incurred by the quantization operation, consistent with the analysis of OpenVLA Kim et al. (2024). Additionally, we accelerated the quantized model using the Spec-VLA framework, discovering that SD could speed up the quantized verification model. It achieves a significant speedup compared with AR decoding (Table 6).
| Dataset | OpenVLA | |
| Precision | Speedup | |
| LIBERO-Goal | bf16 | 1.00 |
| int8 | 0.24 | |
| int4 | 0.61 | |
| LIBERO-Object | bf16 | 1 |
| int8 | 0.21 | |
| int4 | 0.59 | |
| LIBERO-Spatial | bf16 | 1 |
| int8 | 0.23 | |
| int4 | 0.55 | |
| LIBERO-Long | bf16 | 1 |
| int8 | 0.23 | |
| int4 | 0.57 | |
| Dataset | SpecVLA | |
| Precision | Speedup | |
| LIBERO-Goal | bf16 | 1.42 |
| int8 | 1.61 | |
| int4 | 1.34 | |
| LIBERO-Object | bf16 | 1.38 |
| int8 | 1.41 | |
| int4 | 1.33 | |
| LIBERO-Spatial | bf16 | 1.28 |
| int8 | 1.31 | |
| int4 | 1.29 | |
| LIBERO-Long | bf16 | 1.22 |
| int8 | 1.32 | |
| int4 | 1.15 | |